OpenStudy (anonymous):

What is the cubed root of the square root of 64x^6?

OpenStudy (saifoo.khan):

4x^2

OpenStudy (anonymous):

=(8x^3)^1/3=2x

OpenStudy (anonymous):

so which is it lol?

myininaya (myininaya):

\[\sqrt{64*x^6}=8x^3\] \[\sqrt[3]{8x^3}=2x\]

OpenStudy (dumbcow):

2x

OpenStudy (saifoo.khan):

4^3 = 64

OpenStudy (anonymous):

sqrt = 8x^3 cube root of that = 2x.

OpenStudy (anonymous):

Now do it the other way around:-)